# How Animals Dance (When You’re Not Looking)

### v1

@@ 3.3. Choreography Pattern Driven Dance Synthesis
To ensure both expressive motion and synthesis feasibility, we filter nodes based on the average per-pixel flow magnitude $|F(I_{u},I_{v})|$, computed over the foreground region of the subject. Nodes with flow that is too small (insufficient motion) or too large (challenging for video synthesis) are discarded, as shown in Fig. 6 . Only node pairs with acceptable motion range are retained:
(4) $V=\left\{(I_{u},I_{v})\mid{M}_{\text{low}}<|F(I_{u},I_{v})|<{M}_{\text{high}}\right\}$
@@
(5) $E=\left\{\left((I_{u},I_{v})\rightarrow(I_{w},I_{x})\right)\;\middle|\;|F(I_{v},I_{w})|<{M}_{\text{high}}\right\}$
Refer to caption Refer to caption
(a) $|F(I_{u},I_{v})|=9.03$
Refer to caption Refer to caption
(b) $|F(I_{u},I_{v})|=72.52$
Figure 6. Keyframe pairs with excessively small or large motion are excluded from the graph nodes.
@@
Graph optimization. We define a node assignment function:$\phi:\mathcal{L}\rightarrow V$, which maps each choreography label $l_{i}\in\mathcal{L}$ to a graph node $(I_{u},I_{v})\in V$, forming a walk path through the keyframe graph. The goal is to find the optimal assignment $\phi^{*}$ that minimizes the total transition cost across the sequence:
(6) $\phi^{*}=\text{argmin}\quad\sum_{i=0}^{\frac{N}{2}-2}\mathcal{T}(\phi(l_{i}),\phi(l_{i+1}))$
@@
(7) $l_{i}=l_{j}\;\Leftrightarrow\;\phi(l_{i})=\phi(l_{j})$
(8) $l_{i}^{\prime}=l_{j}\;\Leftrightarrow\;\phi(l_{i})=\mu(\phi(l_{j}))$
@@
(10) $l_{i}\neq l_{i+1},\;l^{\prime}_{i}\neq l_{i+1}\;\Rightarrow\;(I_{a}\neq I_{c}\land I_{b}\neq I_{d})$
@@ 3.4. Warp to Music
We generate the final dance video by applying a video diffusion model to synthesize in-between frames along the optimized keyframe walk path $\mathcal{P}=\{I_{p_{0}},I_{p_{1}},\dots,I_{p_{N-1}}\}$, where each keyframe $I_{p_{i}}$ corresponds to beat position $i$. Note that since there are motion repetition in the choreography, we only have to synthesize videos between unique keyframe pairs. In practice, we use Framer (Wang et al., 2024a), which generates $14$ in-between frames, and we assume a fps of $25$. To synchronize with the music, we warp the video timeline such that the timing of every keyframe in $\mathcal{P}$ align with the corresponding beat time in the audio. Following the visual rhythm strategy from (Davis and Agrawala, 2018), we accelerate the warping rate into beat points and decelerate before and after to preserve beat saliency while ensuring temporal smoothness.
4. Experiments
4.1. Keyframes Generation
We generate initial keyframes set by prompting a text-to-image model, such as FLUX to generate an image grid with consistent keyframes using prompt template like “a 3x2 grid of frames, showing 6 consecutive frames from a video clip of [DESCRIPTION PROMPT]”. For example, the description prompt might be “a marmot dancing wildly in the wild alpine landscape, striking a variety of fun and energetic poses.”. The same prompt format can also be used with GPT-4o’s image generation modules, which supports even finer specifications. For instance, one can specify: “Generate a grid with 2 rows and 3 columns. Depict a quokka with distinct poses with a wild background. The postures should feel natural for the quokka’s anatomy….”. In all of our results, we use a total of $6$ input keyframes.
4.2. Choreoraphy Pattern Labeling
We collect a total of $20$ dance video clips featuring various 4/4 music tracks from Youtube and Tiktok, ranging from $12$s to $25$s in length. To create ground truth, we manually annotate the choreography pattern label sequence. We then evaluate our method in Section 3.1 by comparing our extracted label sequences against the ground truth ones. Beat times are detected using Librosa, and SMPL-X pose sequences are recovered from the videos using GVHMR (Shen et al., 2024). We set the threshold values as follows: $\epsilon_{\theta}=0.21,\epsilon_{\theta^{\prime}}=0.25$ and $\epsilon_{d}=0.1$. Specifically, we evaluate two aspects: (1) Clustering accuracy, where each unique label—including mirrored variants—is treated as a distinct cluster. We assess the clustering results using standard metrics: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI); (2) Mirror detection accuracy, where we compute precision and recall based on the correctly identified mirrored motion segments pairs.
Clustering Mirroring
[…]
0.94 0.98 0.93 0.91 0.92
Table 1. Evaluation on choreography pattern labeling.
Figure 7. User study results. Green: Beat accuracy; Red: Dance perception; Purple: Anatomy plausibility; Brown: Visual coherence.
4.3. Dance Results
@@
We generate a collection of input keyframe grids featuring approximately $25$ animal instances. The animal classes include marmot, capybara, hedgehog, meerkat, penguin, sea otter, cat, raccoon, quokka, beaver and others. We also incorporate anthropomorphic characters such as Elmo. For the video results, we generate dance sequences for these instances using six popular song clips, ranging from $16$ to $30$ seconds in length, with choreography patterns extracted from the corresponding YouTube videos. While we showcase our method using these examples, it can adapt to any choreography patterns paired with music. Selected results are shown in Fig. 9 and Fig. 12, and we highly recommend viewing the supplementary video for the full experience.
User study. Since there are no real animal dance videos or existing methods for direct comparison, we conducted a user study to evaluate our approach. We used $40$ generated dance videos across $6$ different songs, and invited $31$ participants to take part. Each participant was presented a random set of $8$ dance videos and asked to rate the following aspects on a scale from 00 to $5$: (1) Beat accuracy—Does the movement of the animal appear synchronized with the music beat? (2)Dance perception—How confident are you that this is a dance rather than random motion? (3) Anatomical plausibility—Does the animal’s appearance align with its anatomical structure? (4) Visual coherence— Does the video looks visually coherent? The responses are shown in Fig. 7. Our generated dance videos obtained average score of $4.37$, $4.35$, $4.11$, and $3.70$ for these factors, respectively.
4.4. Dance Controllability
Motion range control. The threshold parameters $M_{\text{low}}$ and $M_{\text{high}}$ control the range of allowed motion magnitudes between keyframes. A typical setting is $M_{\text{low}}\geq 12.$ and $M_{\text{high}}\leq 60.$. Within this range, lowering $M_{\text{low}}$ and increasing $M_{\text{high}}$ introduces more candidate nodes and transitions, potentially resulting in richer and more expressive dances.
Figure 8. Keyframe pose grid mimicking. Top row: A keyframe pose grid template showing six distinct poses of a capybara. Middle and bottom rows: A meerkat and a hedgehog mimicking the capybara’s poses, while preserving their own anatomical structure and personality, generated using GPT-4o.
@@ C D C’ D’
Figure 9. Selected keyframes from our generated dance video set to the Bumblebee song. The choreography pattern follows a reference dance extracted from Youtube video [GIq7ZgmxE2w], from $15.5$s to $45.5$s: A-B-A’-B’-C-D-C’-D’-A-B-A’-B’-C-D-C’-D’-E-F-G-G-G-G-H-H-H’-H’-G-G-G-G-H-H.
5. Failures
In the keyframe graph, we use average per-pixel flow magnitude between two keyframes as a proxy for motion strength. However, this measure becomes unreliable in cases where the two keyframes represent mirrored side views, where the flow magnitude fails to reflect the true motion complexity between poses. For example, in Fig. 10, the average flow magnitude from the left one to the right one is $|F(I_{u},I_{v})|=36.81$, which appears moderate due to incorrect correspondences between opposite sides. In reality, this motion is difficult for the video model to synthesize.
Refer to caption Refer to caption
Figure 10. Failures. While average flow magnitude from the left one to the right one $|F(I_{u},I_{v})|=36.81$ is not large, the motion is hard to synthesize for video model.
6. Discussions
6.1. Comparison with Anthropomorphic Animation
Another potential relevant line of work for generating animal dance videos is character animation, which animates an input image using a pose sequence, typically extracted from human motion videos. However, transferring human poses to animals remains an open challenge due to the need to solve complex correspondence problems across different body morphologies and modalities, let alone dealing with long and diverse pose sequences as in dance. As an example, we experimented with generating such videos using Animate-X (Tan et al., 2024): we extract pose sequences from a real dance video and use GPT-4o to generate an input image of a marmot that mimics the human pose from the first frame, providing a better start point for animation. As shown in Figure 11, the results looks blurry and the animated marmot taking on human-like body shapes.
@@ 6.1. Comparison with Anthropomorphic Animation
Figure 11. Example results using character animation method. Top row: pose sequence extracted from real dance video (Youtube shorts, [jkRIIH42Vo8]). Bottom row: animation results from applying the pose sequence to the leftmost image using Animate-X (Tan et al., 2024).
6.2. Limitations
We use an offline video diffusion model to generate short motion segments between keyframes, for example, $0.5$s for a $120$ BPM song. While the results are often visually coherent, , the motion can sometimes appear unrealistic; animals may appear to slide or morph between poses rather than moving in a physically plausible way. This reflects the current limitations of video diffusion models in producing naturalistic motion for articulated subjects. However, with continued advances in large-scale video diffusion models, we are optimistic that these issues can be addressed in the future.
A A’ B B’
C D E E’
(a) Selected keyframe walk path set to Uptown Funk song. Choreography pattern extracted from Youtube video [U9Zj1BaH01c], from $16.0$s to $36.0$s: A-A’-A-A’-A-A’-A-A’-B-B’-C-D-E-E’-E-E’.
A A’ B C
D D’ E E
(b) Selected keyframes from our generated dance video set to the APT. song. The choreography pattern follows a reference dance extracted from Youtube video [DJz1zlm73HI], from $11.0$s to $37.0$s: A-A’-A-A’-A-A’-B-C-A-A’-A-A’-A-A’-B-C-D-D’-D-D’-E-E-E’-E’-D-D’-D-D’-E-E-E’-E’
A A’ B
@@
C D D
@@
(c) Selected keyframes from our generated dance video set to the Can’t Stop Feeling song. The choreography pattern follows a reference dance extracted from Youtube video [xyMBnn3dzdU], from $59.0$s to $79.0$s: A-A’-B-C-D-D-D-D-E-E’-E-E’-F-F’-F-F’-A-A’-B-C-D-D-D-D-E-E’-E-E’-F-F’-F-F’.
Figure 12. Dance results. Keyframe pairs are labeled by the choreography pattern label. The sequence should be read from left to right across the top row, followed by left to right across the bottom row. See supplementary for the full dance video with music.
6.3. Future Work
To generate more advanced and musically aligned dances, two directions can be explored: (1) dance motion realism: the motions generated by the video diffusion model may not always reflect plausible or expressive dance motion. Incorporating priors that favor natural, dance-like movement could improve alignment with musical context. (2) style compatibility: although our method follows the choreography pattern, it does not consider musical style. Modeling genre-specific movement characteristics could enhance the stylistic coherence of generated dances.
6.4. Conclusion
We present a novel keyframe-based paradigm for generating music-synchronized, choreography-aware animal dance videos. Our work opens up exciting opportunities for creative and fun applications of dancing animals in entertainment and social media.
Acknowledgments. We thank Tong He and Baback Elmieh for helpful discussions and feedback. This work was supported by the UW Reality Lab and Google.

### v2

@@ 3.3 Choreography Pattern Driven Dance Synthesis
To ensure both expressive motion and synthesis feasibility, we filter nodes based on the average flow magnitude $|F(I_{u},I_{v})|$ from $I_{u}$ to $I_{v}$, computed over the foreground region of the subject. We use RAFT [teed2020raft] to compute the optical flow. Nodes with flow that is too small or too large are discarded. Only node pairs with acceptable motion range are retained:
$V=\left\{(I_{u},I_{v})\mid{M}_{\text{low}}<|F(I_{u},I_{v})|<{M}_{\text{high}}\right\}$ (4)
@@
$E=\left\{\left((I_{u},I_{v})\rightarrow(I_{w},I_{x})\right)\;\middle|\;|F(I_{v},I_{w})|<{M}_{\text{high}}\right\}$ (5)
@@
[…]
$\phi^{*}=\text{argmin}\quad\sum_{i=0}^{\frac{N}{2}-2}\mathcal{T}(\phi(l_{i}),\phi(l_{i+1}))$ (6)
@@
$l_{i}=l_{j}\;\Leftrightarrow\;\phi(l_{i})=\phi(l_{j}),\ l_{i}^{\prime}=l_{j}\;\Leftrightarrow\;\phi(l_{i})=\mu(\phi(l_{j}))$ (7)
@@
$l_{i}\neq l_{i+1},\;l^{\prime}_{i}\neq l_{i+1}\;\Rightarrow\;(I_{a}\neq I_{c}\land I_{b}\neq I_{d})$ (9)
@@ 3.4 Warp to Music
We generate the final dance video by applying a video diffusion model to synthesize in-between frames along the optimized keyframe walk path $\mathcal{P}=\{I_{p_{0}},I_{p_{1}},\dots,I_{p_{N-1}}\}$, where each keyframe $I_{p_{i}}$ corresponds to beat position $i$. Note that since there are motion repetition in the choreography, we only have to synthesize videos between unique keyframe pairs. In practice, we use Framer [wang2024framer], which generates $14$ in-between frames, and we assume a fps of $25$. To synchronize with the music, we warp the video timeline such that the timing of every keyframe in $\mathcal{P}$ align with the corresponding beat time in the audio. Following the visual rhythm strategy from [davis2018visual], we accelerate the warping rate into beat points and decelerate before and after to preserve beat saliency while ensuring temporal smoothness.
4 Experiments
@@
We generate a collection of input keyframe grids featuring approximately $25$ animal instances, some captured as half-body views. The animal classes include marmot, capybara, hedgehog, meerkat, penguin, sea otter, cat, quokka, beaver and others. We also incorporate characters such as Elmo. For the video results, we generate dance videos for these instances using five popular song clips, ranging from $16$ to $28$ seconds in length, with choreography patterns extracted from the corresponding YouTube video clips. While we showcase our method using these examples, it can adapt to any choreography patterns paired with music. Fig. 4 shows selected examples of the final keyframe pairs assigned for each choreography label, arranged in the order specified by the choreography pattern. We highly recommend viewing the supplementary video for the full experience.
A A’ B B’
C D E E’
(a) Dance to Uptown Funk following the choreography pattern extracted from Youtube [U9Zj1BaH01c] ($16.0$s to $36.0$s): A-A’-A-A’-A-A’-A-A’-B-B’-C-D-E-E’-E-E’.
@@ C D C’ D’
(b) Dance to Bumblebee following the choreography pattern extracted from Youtube [GIq7ZgmxE2w] ($15.5$s to $45.5$s): A-B-A’-B’-C-D-C’-D’-A-B-A’-B’-C-D-C’-D’-E-F-G-G-G-G-H-H-H’-H’-G-G-G-G-H-H.
A A’ B C D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) Dance to Rasputin following the choreography pattern extracted from Youtube [jkRIIH42Vo8] ($12.0$s to $33.24$s): A-A’-A-A’-A-A’-A-A’-B-B-B-B-B-B-B-B-C-C-C-C-D-D’-D.
A A’ B C D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(d) Dance to APT following the choreography pattern extracted from Youtube [DJz1zlm73HI] ($11.0$s to $37.0$s): A-A’-A-A’-A-A’-B-C-A-A’-A-A’-A-A’-B-C-D-D’-D-D’-E-E-E’-E’-D-D’-D-D’-E-E-E’-E’.
Figure 4: Selected examples from our generated dances. Keyframe pairs are labeled by the choreography pattern label, arranged in the order specified by the choreography pattern. For clarity, we show only a portion of the full sequence here. See supplementary for the complete dance video with music.
4.1 Keyframes Generation
We generate initial keyframes set by prompting text-to-image model FLUX to generate an image grid with consistent keyframes using prompt template like “a 3x2 grid of frames, showing 6 consecutive frames from a video clip of […]”. For example, the description prompt might be “a marmot dancing wildly in the wild alpine landscape, striking a variety of fun and energetic poses”. The same prompt format can also be used with GPT-4o, which supports even finer specifications. For instance, one can specify: “Generate a grid with 2 rows and 3 columns. Depict a quokka with distinct poses with a wild background. The postures should feel natural for the quokka’s anatomy….”. In all of our results, we use a total of $6$ input keyframes.
4.2 Choreography Pattern Labeling
Clustering Mirroring
[…]
0.94 0.98 0.93 0.91 0.92
Table 1: Evaluation on choreography pattern labeling.
We collect a total of $20$ dance video clips featuring various 4/4 music tracks from Youtube and TikTok, ranging from $12$s to $28$s in length. To create ground truth, we manually annotate the choreography pattern label sequence. We then evaluate our method in Section 3.1 by comparing our extracted label sequences against the ground truth ones. Beat times are detected using Librosa, and SMPL-X pose sequences are recovered from the videos using GVHMR [shen2024gvhmr]. We set the threshold values as follows: $\epsilon_{\theta}=0.21,\epsilon_{\theta^{\prime}}=0.25$ and $\epsilon_{d}=0.1$. Specifically, we evaluate two aspects: (1) Clustering accuracy, where each unique label—including mirrored variants—is treated as a distinct cluster. We assess the clustering results using standard metrics: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI); (2) Mirror detection accuracy, where we compute precision and recall based on the correctly identified mirrored motion segments pairs.
4.3 Baseline Comparisons
Baselines. A straightforward baseline is music-conditioned video generation, which generates videos directly from audio and text prompts. However, methods such as MusicInfuser [hong2025musicinfuser] are trained on human dance videos and typically produce only short clips, e.g., 5s. So they cannot generalize to long animal dance videos synchronized to the input music. Several examples are provided in the supplementary videos.
Next we compare our method to human pose driven single image animation method using Animate-X [tan2024animate], which animates an input image according to a sequence of human skeleton poses and works with anthromorphic characters. For each of our generated dance videos, we extract the pose sequence from the same reference video used to extract the choreography pattern and use it as the driving sequence for Animate-X.
Figure 5: User ratings of our approach compared to Animate-X on various criteria.
User study. Since there are no existing long, structured animal dance videos for direct reference, we conducted a perceptual user study to evaluate the generated dance videos. We used $40$ generated dance video pairs across $5$ different songs, and invited $31$ participants. Each participant was presented a random set of $8$ pairwise comparisons of our results and Animate-X. The order of the videos within each pair was randomized. For each pair, participants were asked to choose which video they judged better on each of four criteria, or select “similar” if they found no difference: (1) Beat accuracy—are the dances synchronized with the music beats? (2) Appearance & motion naturalness —do the animals’ body proportions and movements feel natural for them? (3) Visual quality—do the videos have high overall visual quality (e.g. sharpness, clarity, and fewer artifacts)? (4) Motion fluidity—are the dance movements smooth and fluid? The responses are shown in Fig. 5, and example qualitative comparisons are shown in Fig.6.
Discussions. While beat accuracy was rated similarly for both methods, participants found our results more natural-looking for animals and of higher visual fidelity than those from Animate-X. Specifically, $80.5\%$ of responses rated our videos superior in appearance and motion naturalness, and $88.5\%$ rated them higher in overall visual quality. Animate-X was preferred for motion fluidity ($59.3\%$).
These results are in line with the different setups of these two methods. Animate-X generates animal dances by following fine-grained per-frame human pose sequences, which naturally leads to human-like figures and dance motions—resulting in more fluid and richer movement compared to our method. Yet transferring human poses to animal bodies is inherently difficult: it requires to solve complex correspondence across different body morphologies, and becomes even more challenging when handling the long and diverse pose sequences of real dances. For example, in Fig. 6, Animate-X maps the human arms to the penguin’s wings, causing the wings to move like human arms; the penguin’s differently shaped head further introduces blurry artifacts. Our method instead uses choreography pattern as higher-level control, letting the video diffusion model generate in-between motions so that the final dance follows the intended dance structure rather than fined-grained poses.
Animate-X
Ours
Animate-X
Ours
Figure 6: Sample frames from both our results (cropped for visualization) and Animate-X. Top: APT (Youtube [DJz1zlm73HI]); Bottom: Cann’t Stop Feeling (Youtube [xyMBnn3dzdU]) used as the reference dance videos. The red box marks the input image for Animate-X. See supplementary for the full video comparison.
5 Discussions
5.1 Limitations
We use an offline video diffusion model to generate short motion segments between keyframes (e.g., $0.5$s for a $120$ BPM song). The motion can sometimes look unrealistic: animals may appear to slide or morph between poses rather than moving in a physically plausible way. This stems from the limitations of current video diffusion models in producing naturalistic motion for articulated subjects. However, we are optimistic that these issues can be addressed with continued advances in large-scale video diffusion models.
5.2 Future works
To generate more advanced and musically aligned animal dances, two directions can be explored: (1) dance motion realism: the motions generated by the video diffusion model may not always reflect plausible or expressive dance motion. Incorporating priors that favor natural, dance-like movement could improve alignment with musical context. (2) style compatibility: although our method follows the choreography pattern, it does not consider musical style. Modeling genre-specific movement characteristics could enhance the stylistic coherence of generated dances.
5.3 Conclusions
We present a paradigm for generating music-synchronized, choreography-aware animal dance videos by using choreography pattern as a novel control to impose a structure on the keyframes input. Additional analysis of the advantages of this framework, especially in terms of control, as well as its uncommon failure cases, is included in the accompanying appendix. Overall, our work opens up exciting opportunities for creative and fun applications of dancing animals in entertainment and social media.
Acknowledgments. We thank Tong He, Baback Elmieh, Susung Hong, and Meng-Li Shih for helpful discussions and feedback. This work was supported by the UW Reality Lab and Google.
Please see our supp. video for a full set of video results.
Appendix A User Control
@@
Figure 7: Keyframe pose grid mimicking. Top row: A keyframe pose grid template showing six distinct poses of a capybara. Middle and bottom rows: A meerkat and a hedgehog mimicking the capybara’s poses, while preserving their own body structure, generated using GPT-4o.
Motion range control. The threshold parameters $M_{\text{low}}$ and $M_{\text{high}}$ control the range of allowed motion magnitudes between keyframes. A typical setting is $M_{\text{low}}\geq 12.0$ and $M_{\text{high}}\leq 60.0$ at resolution $1024\times 576$. Within this range, lowering $M_{\text{low}}$ and increasing $M_{\text{high}}$ introduces more candidate nodes and transitions, potentially resulting in richer and more expressive dances.
Appendix B Failure Cases
@@
Figure 8: Background inconsistency caused by missing background Canny edges in the original keyframes. Top row: original keyframes generated by GPT-4o. Second row: Canny edge maps used in generating the refined keyframes. Third row: our refined keyframes. Bottom row: zoom-in of the top-left corners (red rectangles) of the refined keyframes, highlighting slight background inconsistencies where Canny edges are absent.
Background consistency. As described in Section 3.2 in the main paper, we re-generate the keyframes with the fine-tuned model to improve visual consistency in the initial keyframe set using a shared background Canny edge map which is inpainted from the original keyframes as control. However, slight background consistency can remain when the original keyframes have shallow depth of field, which limits the Canny edge extraction. For example, in Fig. 8, the grass background behind the capybara is blurred in the original keyframe, leaving no detectable edges in that region. As a result, the refined keyframes show slight background inconsistencies in the grass (see bottom row). However, for scenes with clear and sharp backgrounds, our method maintains consistency well.
Refer to caption Refer to caption
Figure 9: Failures. The average flow magnitude from left to the right is not large, but the underlying motion intensity is much higher.
Motion intensity estimation. Within the keyframe graph, we estimate the underlying motion strength between keyframe pairs using the average flow magnitude. This measure can be unreliable in certain cases. For example, when two keyframes depict mirrored side views, the flow fails to capture the true motion complexity between poses. In Fig. 9, the average flow magnitude is $36.81$ (image size is $1024\times 576$), which appears moderate due to incorrect correspondences between opposite sides, but the sea otter must rotate from one side view to the other, and this motion is more complex and sometimes challenging for the video model to synthesize. Establishing a reliable link between keyframe flow and the generated motion strength remains an open problem.